OpenStudy (anonymous):

Find the common ratio for the following sequence. 1/2, -1/4, 1/8, -1/16, ...

OpenStudy (phi):

if you have a sequence a,b,c,... find the ratio of b/a (or c/b , etc)

OpenStudy (texaschic101):

to find the common ratio, divide the 2nd number by the first number.. (-1/4) / (1/2) = -1/4 * 2/1 = -2/4 which reduces to -1/2 and as you can see.. 1/2 * -1/2 = -1/4 -1/4 * -1/2 = 1/8 1/8 * -1/2 = - 1/16 so if you multiply -1/2 by the number, you will get the next number. So -1/2 is the common ratio

OpenStudy (mathmale):

More explicitly, determine the multiplier which takes you from your first term (1/2) to your second (-1/4), from (-1/4) to (1/8). If you can find such a number / multiplier, it's your "common ratio."

OpenStudy (anonymous):

Only 99's commented.

OpenStudy (anonymous):

haha